One mole of a monatomic perfect gas initially at temperature T expands from volume V to 2V, (a) at constant temperature, (b) at constant pressure. Calculate the work of expansion and the heat absorbed by the gas in each case? (5marks)


1
Expert's answer
2021-04-14T05:14:18-0400




At constant Temperature


"dw = RT\ln{V_2\over V_1}"


"dw = RT\ln{2V\over V}"


"\boxed{dw = RT\ln{2}}"


from first law of thermodynamics


dQ = du + dw


since temperature is constant , change in internal energy du = 0

"\implies"

dQ= dw


"\therefore" Heat absorbed "\boxed{dQ = RT\ln{2} }"




At constant pressure


The first law of thermodynamic says


"\boxed{Q=\Delta U+W}"

The change in internal energy of one mole of a monoatomic gas


"\Delta U=\frac{3}{2}P(V_2-V_1)\\
=\frac{3}{2}P(2V_0-V_0)=\frac{3}{2}PV_0=\frac{3}{2}RT_0"

The work done by a gas


"\boxed{W=P(V_2-V_0)=PV_0=RT_0}"

Finally


"\boxed{Q=\frac{3}{2}RT_0+RT_0=\frac{5}{2}RT_0}"